OpenStudy (anonymous):

Don't have answer sheet & like confirmation on my calc: x dy/dx sin y/x = y sin y/x – x let y=vx then v = y/x & dy/dx = v + (x)dv/dx x[v + (x)dv/dx] sin(v) = vxsin(v) – x substitute values (vx + (x²)dv/dx) = vx – [x/sin(v)] Divided by sin(v) (x²)dv/dx = - x/sin(v) vx cancel on both sides (1) dv/dx = - 1/[(x)sin(v)] divide by (x²) on both sides sin(v) dv/dx = - 1/x multiply by sin(v) on both sides ∫sin(v) dv = - ∫1/x dx - cos(v) = - lnǀxǀ + C - cos(y/x) = - lnǀxǀ + C cos(y/x) - lnǀxǀ + C = 0

OpenStudy (unklerhaukus):

youve solved the DE correctly i got the same result with the same basic method however i found it easier to rearrange the equation before substitution x dy/dx sin y/x = y sin y/x – x dy/dx sin y/x = y/x sin y/x -1 dy/dx = y/x - 1/sin y/x dy/dx = y/x- csc y/x now subsituting y/x=v v + xdv/dx = v-csc v xdv/dx = - csc v sin v dv = -dx/x